OpenStudy (anonymous):

Evaluate ∫(0,2) ∫(y^2,4) (sqrt(x))sinxdxdy...by using integration by parts

OpenStudy (anonymous):

normally is fine too, without integration by parts. But integration by parts would be more helpful!

OpenStudy (anonymous):

...HMM well that's an odd notation... is it: \[\int^2_0\int^4_{y^2}\sin(x)\sqrt{x}\phantom{.}dx\phantom{.}dy\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ALRIGHT! Well I would first find the primitive of \(\sin(x)\sqrt{x}\) first before doing anything! Just to point out, the integral is not very nice according to this: http://www.wolframalpha.com/input/?i=integral+of+sin%28x%29sqrt%28x%29 BUT I'LL DO MY BEST!

OpenStudy (anonymous):

Ok, by the way. If you are doing it the normal way it is much nicer switching the order. But if not it will be very difficult. But I will get extra credit if I use parts and I really need the extra credit haha.

OpenStudy (anonymous):

Haha I give up...I tried but I think im not getting anywhere... here is what I had so far... but I don't think that's going anywhere...SORRY. \[\eqalign{ &\int\sin(x)\sqrt{x}\phantom{.}dx \\ =&\int \sin(x)u\phantom{.}dx\to\phantom{space}u=\sqrt{x}\to u^2=x\to dx=2u\phantom{.}du\\ =&\int\sin(u^2)(u)(2u)du\\ =&2\int u^2\sin(u^2)du \\ =&2\int a\phantom{.}db \\ &\\ a&=u^2\phantom{...}db=\sin(u^2) \\ da&=2u\phantom{....}b=\small{\int}\sin(u^2) \\ &\\ =&2\int a\phantom{.}db \\ =&2\left[ab-\int b\phantom{.}da\right] \\ =&2\left[u^2\int\sin(u^2)-\int2u\left(\int\sin(u^2)\right)\right] \\ }\]

OpenStudy (anonymous):

its ok thanx though!!

OpenStudy (anonymous):

Could you help me do the integral with by switching the order?

OpenStudy (anonymous):

@zpupster hey, where are you in the problem